OpenStudy (anonymous):

I need to find a polynomial for the perimeter and the area the square is y+5 on the top and y on the side

OpenStudy (anonymous):

perimeter = 2 y + 2 (y + 5)= 4 y + 10 area = y(y + 5) = y^2+5y